OpenStudy (anonymous):

Cerise can paint the living room in 2 hours and Kyle can paint it in 3 hours. If they work together, how long will it take them to complete the job?

OpenStudy (saifoo.khan):

1/2 + 1/3 = 1/x 1/x = 5/6 x = 6/5

OpenStudy (anonymous):

How would that work?

OpenStudy (saifoo.khan):

need more explaination/.

OpenStudy (anonymous):

Yes, please.

OpenStudy (saifoo.khan):

alright, to solve problems like these.. v take their reciprocals..

OpenStudy (anonymous):

Yes, I just don't understand how you got 5/6.

OpenStudy (saifoo.khan):

u have to take the LCM.

OpenStudy (anonymous):

oh yes..

OpenStudy (saifoo.khan):

got that?

OpenStudy (anonymous):

yep. thanks.

OpenStudy (saifoo.khan):

=]

OpenStudy (saifoo.khan):

u r welcome.